### v2

# PankRAG: Enhancing Graph Retrieval via Globally Aware Query Resolution and Dependency-Aware Reranking Mechanism

Abstract
Recent graph-based RAG approaches leverage knowledge graphs by extracting entities from a query to fetch their associated relationships and metadata. However, relying solely on entity extraction often results in the misinterpretation or omission of latent critical information and relationships. This can lead to the retrieval of irrelevant or contradictory content, as well as the exclusion of essential information, thereby increasing hallucination risks and undermining the quality of generated responses. In this paper, we propose PankRAG, a framework designed to capture and resolve the latent relationships within complex queries that prior methods overlook. It achieves this through a synergistic combination of a globally-aware hierarchical resolution pathway and a dependency-aware reranking mechanism. PankRAG first generates a globally aware resolution pathway that captures parallel and progress relationships, guiding LLMs to resolve queries through a hierarchical reasoning path. Additionally, its dependency-aware reranking mechanism utilizes resolved sub-question dependencies to augment and validate the retrieved content of the current unresolved sub-question. Experimental results demonstrate that PankRAG consistently outperforms existing state-of-the-art methods, underscoring its generalizability. The code is available at:https://github.com/Ninggggy/PankRAG.
Index Terms—  Question answering, Retrieval-augmented generation, Query decomposition.
1 Introduction
Large language models (LLMs)  [17, 18]exhibit exceptional performance across diverse tasks by leveraging pre-training [11] on massive text corpora to generate appropriate responses. While LLMs’ capabilities can be further enhanced through fine-tuning [2] on downstream tasks, finite parameter capacity limits their ability to store comprehensive and evolving world knowledge. Retrieval-augmented generation (RAG)  [15, 9, 6, 20, 5, 12, 7, 14]has emerged as the predominant solution, owing to its straightforward implementation. It enables static models to access up-to-date external knowledge while preserving interpretability and provenance tracking.
RAG dynamically integrate heterogeneous knowledge bases [21, 3, 8, 16, 13] to enhance contextual reasoning and factual grounding. Among these, knowledge graphs  [1] serve as structured frameworks that encode information through interconnected entities and relationships, explicitly modeling semantic dependencies to enable human-like understanding and multi-hop logical inference. Recently, many approaches have integrated RAG with knowledge graphs to enhance factual grounding and multi-hop inference, with prominent examples including GraphRAG[3], LightRAG[8], HippoRAG[13], and G-Retriever[10]. For instance, GraphRAG[3] enhances question answering by adaptively synthesizing precomputed summaries of entity communities. LightRAG[8] leverages a hierarchical retrieval system that fuses graph-based reasoning with vector matching to ensure efficient knowledge extraction and semantic fidelity for multi-hop queries. However, the single-phase retrieval approach of current graph-based RAG methods is ill-suited for deep reasoning, causing them to misinterpret or omit latent critical information and relationships when handling complex queries. For Specific Complex Queries (SCQ) [8] requiring precise multi-hop inference, they often retrieve broad, irrelevant information centered on a core entity instead of the specific relational facts, increasing the risk of LLMs’ hallucinations. Similarly, for Abstract Complex Queries (ACQ) [8] demanding conceptual synthesis, they tend to fetch superficial definitions while overlooking the rich, implicit context required for a nuanced response.
In this paper, we propose PankRAG, which resolves complex queries through a hierarchical reasoning pathway with global awareness, augmented by a dependency-aware reranking mechanism. Specifically, our approach leverages the robust planning capabilities of LLMs to generate a globally-aware (i.e., holistically planned rather than stepwise) resolution pathway, explicitly revealing parallel and progress relationships among sub-questions. Next, we refine complex queries by integrating the implicit critical information resolved from prior sub-question resolutions. This dual-level process produces a hierarchical resolution pathway with global awareness, significantly enhancing retrieval performance and response quality. It enables LLMs to dynamically disambiguate and contextualize ambiguous or polysemous queries before final answer generation. During execution, our dependency-aware reranking mechanism augments and validates retrieved content for sub-questions. This process refines the retrieved context by integrating resolved sub-question dependencies, thereby enhancing the alignment between retrieved contexts and resolved sub-question dependencies.
Our contributions to this work are as follows:
  • We propose a hierarchical query decomposition strategy that generates a globally-aware resolution pathway, enabling LLMs to capture and resolve latent information and relationships, ensuring a tight alignment between generated responses and query intent.
  • We propose a dependency-aware reranking mechanism that augments and validates retrieved content for the current sub-question by integrating resolved sub-question dependencies.
  • PankRAG demonstrates consistent superiority over state-of-the-art baselines, achieving comprehensive improvements on four ACQ datasets evaluated across six dimensions and two SCQ datasets assessed via four metrics.
Fig. 1: The overall workflow of PankRAG.
2 Methodology
2.1 Graph Contruction
The source document corpus $D$ is segmented into $n$ text chunks $\{D_{i}\}^{n}_{i=1}$. As the Equation 1, for each chunk $D_{i}$, the LLM performs entity recognition and relationship extraction, abstracting entities $V_{i}$ and relations $E_{i}$ from unstructured text via prompting strategies. A knowledge graph $G$ is then constructed by instantiating nodes representing entities $\{V_{i}\}$ and edges representing relationships $\{E_{i}\}$ in Equation 2.
$(\mathcal{V}_{i},\mathcal{E}_{i})=\bigcup_{D_{i}\in\mathcal{D}}\textit{Extract}(D_{i})$ (1)
$\mathcal{G}=\left\{\mathcal{V}_{i},\mathcal{E}_{i}\right\}_{i=1}^{n}$ (2)
Next, Multiple communities $\{Com_{i}\}$ are aggregated into a higher-level community $H_{i}$ via Leiden clustering, and this process is iteratively applied to hierarchically construct a multi-layered community structure. Finally, multi‑scale community summarization is executed by the LLM.
$\mathcal{H}_{i}=\textit{Leiden}((\textit{Com}_{i})_{i=1}^{k})$ (3)
2.2 Globally-Aware Planning
Fig. 2: Operation example of Globally-Aware Planning. [A1.1], [A1.2], and [A2.1] correspond to the resolved outputs of SubQ[1.1], SubQ[1.2], and SubQ[2.1], respectively.
We propose a hierarchical query decomposition strategy that synergizes bottom-up relationship decomposition with top-down contextual rephrasing to address both implicit parallel or progress relationships and ambiguous or polysemous expressions in complex queries. Unlike conventional iterative approaches (e.g., Chain-of-Thought, CoT), which suffer from fixed iteration counts and myopic decomposition prone to local optima, our strategy dynamically adjusts the granularity and sequence of reasoning steps based on the query’s inherent structure.
  • Bottom-up phase: In the Parallel Relationship Detection step, parallel semantic relationships—such as independent sub‑questions or co‑occurring entities—are identified to enable parallel execution, thereby optimizing computational efficiency by minimizing resource contention. If no parallelism is detected, a single resolution sequence is initialized. When parallelism exists, multiple concurrent sequences (SubQ1.x and SubQ2.x in Figure 2) are adaptively spawned to optimize concurrent execution. Concurrently, latent progress relationships are extracted from the query. Upon identifying distinct relationships within the query, sub-questions are generated and formalized into a directed acyclic graph (DAG), where nodes represent sub‑questions and edges encode their relationships. Finally, topological sorting is applied to enforce prerequisite‑resolution order, ensuring logical coherence during execution.
  • Top-down phase: For queries containing explicit ambiguities or polysemous expressions, a disambiguation sequence combining resolved sub-question dependencies (SubQ0.x in Figure 2) is invoked. In this phase, an LLM-driven rephrasing process converts ambiguous queries into logically partitioned formulations.
This dual mechanism reduces the risk of deviation in reasoning paths and enhances computational efficiency by eliminating redundant iterations, achieving a balanced trade‑off between precision and scalability. The detailed workflow is above inFigure 1, and a real-world query in the Multihop-RAG [22] dataset example is presented in Figure 2.
2.3 Execution
Following GraphRAG[3], we tailor our retrieval strategy to the query type. For SCQs, we employ a local search that targets query-referenced nodes and expands to their neighbors, ensuring precise entity-centric grounding. For ACQs, we utilize hierarchical community retrieval, dynamically selecting between detailed leaf-level and thematic high-level summaries to balance retrieval breadth and depth.
Table 1: Overall Performance on the ACQ Datasets.
Agriculture Legal CS Mix
NaiveRAG PankRAG NaiveRAG PankRAG NaiveRAG PankRAG NaiveRAG PankRAG
Comprehensiveness 18.0% 82.0% 9.6% 90.4% 20.4% 79.6% 16.8% 83.2%
Diversity 19.6% 80.4% 7.2% 92.8% 27.6% 72.4% 8.8% 91.2%
logicality 2.8% 97.2% 8.8% 91.2% 7.6% 92.4% 17.2% 82.8%
relevance 6.0% 94.0% 11.2% 88.8% 12.8% 87.2% 19.2% 80.8%
coherence 2.8% 97.2% 8.4% 91.6% 6.8% 93.2% 16.8% 83.2%
Empowerment 12.0% 86.0% 9.6% 90.4% 16.4% 83.6% 11.6% 88.4%
GraphRAG PankRAG GraphRAG PankRAG GraphRAG PankRAG GraphRAG PankRAG
Comprehensiveness 40.8% 59.2% 44.4% 55.6% 48.4% 53.2% 44.8% 55.2%
Diversity 44.8% 55.2% 39.2% 60.8% 39.6% 60.4% 43.6% 56.4%
logicality 32.8% 67.2% 43.6% 56.4% 46.4% 53.6% 30.0% 70.0%
relevance 40.4% 59.6% 38.0% 62.0% 39.6% 60.4% 44.4% 55.6%
coherence 34.4% 65.6% 38.0% 62.0% 42.4% 57.6% 31.2% 68.8%
Empowerment 39.6% 60.4% 36.4% 63.6% 44.0% 56.0% 40.0% 60.0%
LightRAG PankRAG LightRAG PankRAG LightRAG PankRAG LightRAG PankRAG
Comprehensiveness 44.8% 55.2% 48.4% 53.2% 35.0% 65.0% 45.2% 54.8%
Diversity 41.6% 58.4% 40.8% 71.6% 28.0% 72.0% 37.2% 62.8%
logicality 16.4% 83.6% 38.0% 57.2% 41.2% 58.8% 42.8% 57.6%
relevance 35.2% 64.8% 38.0% 56.0% 43.6% 56.4% 43.6% 56.4%
coherence 16.8% 83.2% 38.0% 55.6% 42.0% 58.0% 40.0% 60.0%
Empowerment 44.8% 55.2% 45.2% 58.4% 34.0% 60.0% 36.0% 64.0%
HippoRAG PankRAG HippoRAG PankRAG HippoRAG PankRAG HippoRAG PankRAG
Comprehensiveness 45.6% 54.4% 48.0% 52.0% 42.8% 57.2% 46.8% 53.2%
Diversity 42.8% 57.2% 43.2% 56.8% 36.0% 74.0% 44.0% 56.0%
logicality 31.2% 68.8% 43.6% 56.4% 42.8% 57.2% 44.4% 55.6%
relevance 40.4% 59.6% 41.2% 58.8% 47.2% 52.8% 44.8% 55.2%
coherence 32.4% 67.6% 42.4% 57.6% 39.6% 60.4% 45.2% 54.8%
Empowerment 46.0% 54.0% 46.8% 53.2% 40.8% 59.2% 41.2% 58.8%
2.3.1 Dependency-Aware Reranking Mechanism
The dependency‑aware reranking mechanism is triggered when resolving a sub‑question that depends on previously resolved sub‑questions. It dynamically adjusts retrieval priorities by indirectly leveraging these resolved dependencies. Unlike conventional methods that directly incorporate resolved sub-question dependencies, risking cumulative hallucination errors across iterative reasoning steps, our mechanism mitigates such propagation by reinforcing contextually relevant content and filtering out incongruent evidence through semantic similarity validation. If no dependencies are present, the generation proceeds without reranking, preserving computational efficiency while maintaining robustness against cascading inaccuracies.
The reranking mechanism computes a composite score by weighted summation of two components: the intrinsic quality score $R_{i}$ of retrieved content and the semantic similarity $M_{i}$ between retrieved content and resolved dependent sub-question answers.
$\mathcal{M}_{i}=\textit{Cos}(\textit{ori}_{i}+\textit{sub}_{i})$ (4)
The intrinsic quality score $R_{i}$ of retrieved content is generated and retained during the original retrieval phase [3]. The semantic similarity $M_{i}$ is computed as the cosine similarity between resolved sub-question dependencies $\{sub_{i}\}$ and the initially retrieved text chunks $\{ori_{i}\}$. Each chunk’s similarity score is updated and then linearly combined with its original retrieval quality score to optimize prioritization. The weight parameters $\alpha$ (quality) and $\beta$ (relevance) were selected by testing on the dataset(subsection 3.4), ensuring maximal retrieval precision while mitigating hallucination propagation. This process can be formally described as follows.
$\text{Combined Score}=\alpha\times\mathcal{R}_{i}+\beta\times\mathcal{M}_{i}$ (5)
2.3.2 Generation
$Ans=\textit{LLM}(Q,(D,K))$ (6)
The generation process integrates resolved dependencies $D$ and dependency-enhanced reranking knowledge $K$ into a consolidated context, which is utilized by the LLM to generate reliable responses.
3 Experiment
3.1 Experimental Setup
Table 2: Overall Performance on the SCQ Datasets
Multihop-RAG MuSiQue Average
Method Answer Relevance Faithfulness Context Recall Context Precision Answer Relevance Faithfulness Context Recall Context Precision Answer Relevance Faithfulness Context Recall Context Precision
NaiveRAG 66.73% 73.99% 55.17% 54.00% 66.70% 62.50% 59.41% 58.21% 66.72% 66.70% 57.29% 56.11%
GraphRAG 69.56% 86.80% 62.06% 61.18% 68.94% 69.73% 61.57% 60.40% 69.25% 78.27% 61.82% 60.70%
LightRAG 72.40% 86.78% 65.18% 67.00% 67.92% 68.68% 60.99% 58.42% 68.66% 77.73% 63.08% 62.71%
HippoRAG 72.81% 89.84% 73.69% 72.35% 69.18% 75.68% 62.46% 62.19% 70.98% 82.76% 68.01% 67.27%
PankRAG 74.79% 96.53% 87.89% 84.00% 70.62% 88.75% 64.99% 65.35% 72.71% 92.64% 76.44% 74.68%
  • Datasets We follow the standardized preprocessing pipelines from GraphRAG[3] and LightRAG[8] and evaluate our method on six real‑world datasets, encompassing both SCQ and ACQ. We use the Musique [23] and MultiHop-RAG [22] datasets for SCQ. For ACQ, We evaluate our method on the UltraDomin dataset [19], selecting the four domains used (Agriculture, Legal, Computer Science, and Mix)
  • Baselines We compare PankRAG with the three state-of-the-art methods, including NaiveRAG[20], GraphRAG[3], LightRAG[8] and HippoRAG[13] across all datasets.
  • Evaluation Metrics On the SCQ datasets, we employ four metrics from the Ragas framework [4] to assess our method and the baselines: Answer Relevance, Faithfulness, Context Recall, and Context Precision. Lacking ground-truth for the ACQ dataset, we employ GPT-4o-mini as an automated judge. We conduct blinded, pairwise comparisons between Our evaluation involves calculating win rates for our method and baselines across six dimensions. To mitigate positional bias, the presentation order of answers is alternated, and we report the averaged results. We employed GPT-4o-mini for all core tasks. For text embeddings, we utilized the BAAI/bge-m3 model.
Fig. 3: Impact of a pair of hyperparameters in PankRAG on the Mix(left) and Agriculture(right) dataset
3.2 Overall Performance on the ACQ Datasets
As demonstrated in Table 1, PankRAG consistently surpasses baseline methods across all evaluation metrics and datasets.
PankRAG demonstrates comprehensive superiority, achieving high average win rates across diverse datasets, ranging from 64% to 74%. This dominance extends across all evaluation dimensions, excelling particularly in Relevance (69.8%) and Empowerment (69.7%). This consistent outperformance is a direct result of its core mechanisms—globally-aware pathway resolution and dependency-aware reranking—validating its effectiveness on ACQ.
Fig. 4: Impact of a pair of hyperparameters in PankRAG on the Multihop-RAG dataset[22]
Table 3: Ablation study of PankRAG on ACQ datasets.
Multihop-RAG MuSiQue Average
Method Answer Relevance Faithfulness Context Recall Context Precision Answer Relevance Faithfulness Context Precision Context Recall Answer Relevance Faithfulness Context Recall Context Precision
PankRAG-Pank 67.96% 88.21% 62.94 62.21 67.45% 75.16% 59.90% 58.73% 67.70% 81.69% 61.42% 60.47%
PankRAG-Ank 70.21% 94.45 84.31 67.65 69.02% 81.13% 62.27% 61.23% 69.61% 87.79% 73.29% 68.63%
PankRAG 74.79% 96.53% 87.89% 84.00% 70.62% 88.75% 64.99% 65.35% 72.71% 92.64% 76.44% 74.68%
3.3 Overall Performance on the SCQ Datasets
As demonstrated in Table 2, PankRAG consistently achieves the best performance across all experimental settings. PankRAG establishes its superiority by outperforming baselines by 11-17% on average. This advantage is driven by substantial gains across key metrics, including Faithfulness (10-26%), Context Precision(7-18%), Context Recall (8-19%), and Answer Relevance(2-6%), underscoring its enhanced efficacy in resolving SCQs.
Notably, GraphRAG[3] and LightRAG[8] show minimal improvements in Context Recall and Context Precision (4%-6% over NaiveRAG), indicating negligible enhancement in retrieved context quality. This aligns with the identified challenge that LLMs struggle to interpret hidden relationships or information in SCQ, leading to misaligned retrieval targets. In contrast, PankRAG’s significant improvements in Context Recall and Context Precision suggest effective mitigation of this issue, attributed to its globally-aware pathway resolution generation module and dependency-aware reranking mechanism. It enables LLMs to better capture hidden relationships and thereby improve context relevance. Concurrent enhancements in Faithfulness and Answer Relevance further indicate reduced hallucination risks and higher answer quality, underscoring PankRAG’s comprehensive superiority over existing graph-based RAG methods.
Table 4: Ablation study of PankRAG on ACQ datasets.
Agriculture Legal CS Mix
w/o Pank w/o Ank w/o Pank w/o Ank w/o Pank w/o Ank w/o Pank w/o Ank
Comprehensiveness 46.8% 53.2% 34.8% 55.2% 47.2% 52.8% 46.0% 54.0%
Diversity 42.3% 58.8% 38.0% 62.0% 36.2% 54.8% 33.6% 56.4%
logicality 44.4% 55.6% 44.0% 56.0% 44.8% 55.2% 45.2% 54.8%
relevance 45.2% 54.8% 44.8% 55.2% 46.4% 53.6% 46.0% 54.0%
coherence 46.8% 53.2% 43.8% 7.2% 46.4% 53.6% 45.2% 54.8%
Empowerment 56.0% 54.0% 41.2% 58.8% 46.0% 54.0% 43.6% 54.4%
w/o Ank PankRAG w/o Ank PankRAG w/o Ank PankRAG w/o Ank PankRAG
Comprehensiveness 41.6% 58.4% 46.0% 54.0% 46.0% 54.0% 39.2% 55.2%
Diversity 46.4% 53.6% 44.8% 55.2% 41.2% 58.8% 30.8% 56.4%
logicality 33.2% 66.8% 43.2% 56.8% 44.8% 55.2% 32.4% 68.8%
relevance 41.6% 58.4% 46.0% 54.0% 45.6% 54.4% 32.4% 56.4%
coherence 36.8% 63.2% 43.0% 57.0% 43.6% 56.4% 32.4% 60.0%
Empowerment 37.2% 62.8% 43.4% 55.6% 44.0% 56.0% 42.4% 58.8%
3.4 Hyperparameter Analysis
The parameters $\alpha$ and $\beta$ (constrained by $\alpha$ + $\beta$ = 1 to ensure the normalized score range) balance the original quality score against a novel similarity metric derived from resolved dependencies. As shown in Figure 3 and Figure 4, Experimental results demonstrate that the optimal parameters for SCQ are $\alpha$ >0.6 $\beta$ <0.4, whereas for ACQ, $\alpha$ = 0.75 and $\beta$ = 0.25 , with both configurations maximizing retrieval performance.
For SCQ, the optimal performance is achieved with $\alpha$ = 0.6 and $\beta$ = 0.4. This configuration represents a critical trade-off. A higher weight on $\alpha$ ($\alpha$ >0.6) results in insufficient reinforcement from sub-question answers, while a lower weight ($\alpha$ <0.6) leads to an over-emphasis that introduces negative hallucinations, degrading performance. The 0.6:0.4 ratio, therefore, strikes the ideal balance between leveraging sub-questions for context validation and mitigating the risk of performance degradation, yielding the best overall retrieval results. For ACQ, our experiments identified a different optimal configuration of $\alpha$ = 0.75 and $\beta$ = 0.25. This setting was determined as it consistently maximized retrieval performance and achieved the highest win rates against baseline models.
3.5 Ablation Studies
This section’s ablation studies validate the hierarchical contributions of our two core components: globally-aware pathway resolution (Path) and dependency-aware reranking (Ank).
As shown in Table 3 and Table 4, the results demonstrate that the Path module is a foundational component crucial for enhancing retrieval quality. The model variant equipped with Path (PankRAG-Ank) achieves significant improvements in Context Precision and Context Recall over the variant without it (PankRAG-Pank). This finding underscores the necessity of the globally-aware pathway resolution for establishing a high-quality context. Furthermore, the Ank mechanism provides a critical, synergistic enhancement on top of the Path framework. The full PankRAG system, which incorporates the Ank module, shows consistent superiority over the PankRAG-Ank variant across all evaluated metrics. This highlights the critical role of the dependency-aware reranking mechanism in optimizing and refining the final results.
4 Conclusion
In this paper, we propose PankRAG, a novel RAG framework designed to enhance LLMs’ capacity to accurately interpret and retrieve information for complex queries. PankRAG combines a globally aware pathway resolution module with a dependency‐aware reranking mechanism. By executing a globally coherent reasoning pathway and augmenting retrieved content through dependency‐aware reranking, our framework captures and resolves hidden relationships and information in both SCQ and ACQ, significantly enhancing retrieval performance and response quality.